OpenStudy (kyanthedoodle):

Ok. Can someone please teach me how to divide numbers written in scientific notation?

OpenStudy (compassionate):

Subtract the power of the denominator from the numerator

OpenStudy (kyanthedoodle):

I'm still confused.

zepdrix (zepdrix):

Ya :) Deal with the powers of 10 separately. Example:\[\large\rm \frac{2.34\times10^6}{1.52\times10^2}=\frac{2.34}{1.52}\times\frac{10^6}{10^2}\]And then apply exponent rule to the powers of 10, as Comp is describing.

OpenStudy (compassionate):

Say I have \[\frac{ 4 \times 10^3 }{ 2 \times 10^2}\] I would want to divide the numbers like regular and subtract the powers. So I would say 4 divided by 2, and 3 - 2

OpenStudy (hyperlunatic):

I am not needed here

OpenStudy (compassionate):

\[2 \times 10 ^1\] and that would be my solution

OpenStudy (compassionate):

Get it?

OpenStudy (kyanthedoodle):

Ok. Thanks.